OpenStudy (anonymous):

what makes this statement true 9 exponent -2 a 81 b 18 c 1/81 d 1/-81

OpenStudy (raden):

use the property of exponent : a^(-m) = 1/a^m

OpenStudy (raden):

so, 9^-2 = 1/(9^2) = 1/81

OpenStudy (anonymous):

thank you

OpenStudy (raden):

you're welcome